the width of a rectangle is 3 inches less than its length. The area of the rectangle is 340 square inches. What are the length and widthof the rectangle?
Answer by ankor@dixie-net.com(22740)

he width of a rectangle is 3 inches less than its length.
The area of the rectangle is 340 square inches.
What are the length and width of the rectangle?
:
We know the Area is: L * W = 340
Given that,"The width of a rectangle is 3 inches less than its length."
W = L - 3
Replace W with (L-3) in the area equation
L * (L-3) = 340
L^2 - 3L = 340
L^2 - 3L - 340 = 0; a quadratic equation
factors to
(L-20)(L+17) = 0
The positive solution
L = 20 inches is the Length
then
w = 20 - 3 = 17 inches is the width
:
:
Check by finding the area: 20 * 17 = 340